5. Two 1.00 g blocks absorb 1.00 J of heat each. One block is made of gold (Au), and the other one of silver (Ag). How are the changes in their temperature ∆T(Au) and ∆T(Ag) related, and why? Cs(Au) = 0.128 J/(g·◦C), Cs(Ag) = 0.235 J/(g·◦C).

(A) ∆T(Au) = 1.84∆T(Ag) because Cs(Ag) = 1.84Cs(Au)
(B) ∆T(Au) = 1.84∆T(Ag) because Cs(Au) = 1.84Cs(Ag)
(C) ∆T(Ag) = 1.84∆T(Au) because Cs(Ag) = 1.84Cs(Au)
(D) ∆T(Ag) = 1.84∆T(Au) because Cs(Au) = 1.84Cs(Ag)
(E) ∆T(Au) = ∆T(Ag) because of the same masses and amounts of heat
Answer #1

We know that heat absorbed, q = mC×∆T which gives

∆T= q/(mC). So for same value of q and m we have ∆T inversely proportional to C(heat capacity).

Now Cs(Ag)/Cs(Au) = 0.235/0.128 = 1.84. So change in temperature of Au will be 1.84 times change in temperature of Ag.

Thus option (A) is the right answer.